OpenStudy (anonymous):

Suppose that the population of deer in a state is 1,500 and is growing 2% each year. Predict the population after 4 years. A. about 12,000 deer B. about 1,624 deer C. about 3,110 deer D. about 24,000 deer

OpenStudy (anonymous):

\[x_t = x_0(1 + r)^t \\ x_t = 1500(1 + 0.02)^4 \\ x_t = 1500(1.02)^4 \\ x_t = 1500(1.08243216) \\ x_t = \approx 1624\]

OpenStudy (anonymous):

Thanks